Bill is 1.5 times Jane's age. In 7 years, the sum of their ages is 64. What is Bill's age now? Use substitution method. Available answers:
30
20
27
37

Answer by stanbon(75887)

Bill is 1.5 times Jane's age. In 7 years, the sum of their ages is 64. What is Bill's age now? Use substitution method.
z-------
Equations:
b = 1.5j
b+7 + j+7 = 64
-----
Substitute for "b" and solve for "j":
1.5j + 7 + j+7 = 64
2.5j + 14 = 64
2.5j = 50
j = 20 yrs (Jane's age now)
b = 1.5j = 30 years (Bill's age now)
============
Cheers,
Stan H.
=============
Available answers:
30
20
27
37